# What Can RL Bring to VLA Generalization? An Empirical Study

### v2

@@ 1 Introduction
To answer this, we center our investigations on the representative pick-and-place task. On this task, we first evaluate mainstream RL algorithms for large-scale models (PPO (Schulman et al., 2017; Ouyang et al., 2022), DPO (Rafailov et al., 2023; Zhang et al., 2024b), GRPO (Shao et al., 2024; Guo et al., 2025a)) to identify effective VLA fine-tuning strategies. We then conduct a broad comparative evaluation, also on pick-and-place, pinpointing where RL fine-tuning outperforms SFT. As previewed in Fig. 1, this comprehensive evaluation rigorously examines generalization across three key dimensions: 1) Vision: challenging generalization with novel backgrounds (unseen tables) and by overlaying unseen textures on the foreground or the entire image. 2) Semantics: testing understanding via unseen objects, novel receptacles, and varied instruction phrasings to probe language sensitivity and object recognition. 3) Execution: probing robustness by varying initial robot states, object/receptacle positions, and introducing dynamic disturbances like random object reposition within episodes.
@@ 3.2 Vision-Language-Action models
We base our study on OpenVLA (Kim et al., 2024) (Fig. 2), an open-source model that achieves state-of-the-art performance by pairing a fused visual encoder of SigLIP (Zhai et al., 2023) and DINOv2 (Oquab et al., 2023) with a Llama-2 7B language backbone (Touvron et al., 2023), built on the Prismatic VLM (Karamcheti et al., 2024). At each time step the policy receives a single RGB image $o_{t}$ and an instruction $l$, i.e., the history length $H=1$. The image is embedded into visual tokens, the instruction is tokenised with Llama 2’s tokenizer, and the resulting token sequence is fed to the causal transformer decoder.
@@ 4 Effective RL fine-tuning of VLA models
We begin our study by exploring how to fine-tune VLA models with reinforcement learning. As these models inherit a similar scale and structure from their pre-trained LLM/VLM backbones, we first test whether common RL algorithms for large-scale models like LLMs/VLMs transfer effectively to VLAs. In Sec. 4.1, we will show that Proximal Policy Optimization (PPO) (Schulman et al., 2017) consistently improves performance, whereas Direct Preference Optimization (DPO) (Rafailov et al., 2023; Zhang et al., 2024b) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024) struggle to learn something for POMDP robotic tasks. Guided by this finding, we then ablate key design choices within PPO to distill an effective fine-tuning recipe for VLA models in Sec. 4.2.
@@ 4.1 RL algorithms: PPO, GRPO, DPO
  • GRPO estimates the baseline using a group of samples, allowing for direct computation of advantages without explicit value estimation. Although there are differences between natural language and robotic multi-step MDP tasks, we also include a modified version, GRPO(s), where a sampled group of trajectories starts from the same initial state, as suggested by the GRPO paper (Shao et al., 2024). In both settings, we use a batch of trajectories collected from all parallel environments for one task episode as a group.
  • DPO exploits offline datasets annotated with pairwise preferences. In robotic tasks, however, obtaining contrastive trajectories from the same initial state is difficult. Following (Zhang et al., 2024b), we infer trajectory-level preferences from reward signals. To ensure a fair comparison, we employ only the stage-based sparse reward described in Sec. 5.1, rather than the richer reward scheme used in the original implementation by (Zhang et al., 2024b).
These algorithms are illustrated in Fig. 3(a), with further details and implementation specifics provided in Sec. A.1. We evaluate the performance on the pick-and-place task described in Sec. 5.1.
The results are presented in Fig. 3(b), with each experiment conducted using two different random seeds. Our findings indicate that PPO consistently outperforms GRPO. We attribute this to fundamental differences in environment dynamics between natural language tasks and robotic tasks (Li et al., 2024b). We hypothesis that, in robotic tasks’ POMDP, each action sequentially alters the environment state in a non-stationary manner, which may destabilizes GRPO’s advantage estimates. Additionally, PPO surpasses DPO performance. We hypothesize that this is due to the sparse reward structure, which makes it challenging to distinguish between the quality of different trajectories, as well as significant distribution shifts between the offline dataset and interactive execution (Prudencio et al., 2023).
@@ 4.2 Design factors of PPO
Since PPO—and most modern RL methods—use an actor–critic formulation (Konda and Tsitsiklis, 1999), we treat the pretrained VLA policy as the actor and attach a lightweight critic that estimates the state value $V(s)$. To keep the architecture compact, the actor and critic share the entire Transformer backbone; a three-layer MLP value head (see Fig. 4(a)) takes the hidden vector $h^{0}$ produced by the final Transformer block (Vaswani et al., 2017) at the first action-token position and regresses it to a scalar value. To verify its efficacy, we evaluated several critic designs. Feeding the value head with the first action-token embedding $h^{0}$ produced the highest and most stable returns, outperforming the last-token input $h^{n}$ and the concatenation $[h^{0},\dots,h^{n}]$ (see Fig. 4(b)). A critic that uses its own Transformer backbone achieved comparable rewards but trained $35\,\%$ slower and consumed $83\,\%$ more VRAM ($81.3$ GB vs. $44.4$ GB). These results confirm that a shared backbone with $h^{0}$ is the most efficient choice.
@@
We run all experiments with the official OpenVLA checkpoint (Kim et al., 2024) pretrained on the OXE dataset (Collaboration et al., 2023). Due to the poor performance of this checkpoint on our benchmark, we warm it up with $140$ demonstration trajectories gathered by Octo-Small (Team et al., 2024) and a motion planner (see Sec. A.2). As shown in Fig. 5(a), the warm-up model reaches convergence with roughly $50\,\%$ fewer environment steps, while both initialisations attain comparable asymptotic returns given sufficient interaction. Unless stated otherwise, all subsequent experiments use the warmed-up OpenVLA model for initialization.
@@
We show that the update–to–data ratio is a key factor for efficient fine-tuning. In PPO this ratio is set by the epoch hyper-parameter, which specifies how many gradient passes each batch receives. Fig. 5(b) and Fig. 5(c) reveal that increasing the PPO epoch beyond one yields no gains in return or sample efficiency, yet lengthens wall-clock time almost linearly. Accordingly, we fix $\text{epoch}=1$ in all remaining experiments, achieving the fastest training without sacrificing performance.
@@ 5.1 Environments and datasets
To thoroughly investigate the generalization capabilities of VLA, we focus on a typical pick-and-place task (Kim et al., 2025; Li et al., 2024a), where the agent is instructed to place an object from the table into a receptacle. Inspired by prior works (Fan et al., 2025; Stone et al., 2023) and the concept of Vision-Language-Action models, we define three dimensions of generalization: Vision, Semantics (Language), and Execution (Action), as illustrated in Fig. 1.
@@
To probe generalisation, we randomise each task along three axes during training: Vision (16 tables), Semantics (16 objects), and Execution (perturbations of object and receptacle poses). At test time we hold at least one of these factors out of distribution, introducing nine novel objects, sixteen unseen receptacles, five new table surroudings, and sixteen distractor textures. Assets are drawn from Objaverse (Deitke et al., 2023) and other public sources; additional table appearances are synthesised with Stable Diffusion (Rombach et al., 2022) and ControlNet (Zhang et al., 2023a), ensuring surface variation without changing the table’s global pose. More details of asset acquisition and task specification are provided in Secs. A.3 and A.4. Built upon these assets, our tasks run in ManiSkill (Tao et al., 2024) with an 8-DoF WidowX-250S arm. At every step the agent observes a $640\times 480$ RGB frame and a natural-language instruction, and outputs a Cartesian end-effector delta plus a binary gripper signal. Rewards are sparse: $0.1$ for grasping and continuously holding the correct object, and $1.0$ for placing it successfully. For supervised fine-tuning we collect demonstration trajectories using the MPLib motion planner (Guo et al., 2024) and fine-tuned using LoRA (Hu et al., 2022); further dataset details can be found in Sec. A.2.
@@ 5.2 Effect of data scale on SFT performance
Following Lin et al. (2025), we study how supervised fine-tuning (SFT) scales with demonstration count. We train OpenVLA to convergence on datasets ranging from a few hundred to $64$k expert trajectories ($\sim 1.26$M transitions) and report average scores over three random seeds in-distribution and on unseen objects/tables (Fig. 6(a) and 6(b)). We observe that performance plateaus at roughly 16k trajectories in both settings, so we adopt the 16k-trajectory SFT checkpoint as the baseline for comparing RL fine-tuning methods.
@@ 5.3 Performance comparison between RL and SFT
Building on preceding analyses, we plot success rates during RL training for both the training distribution and the OOD object/table split in Fig. 6(c) and 6(d), alongside SFT scores at various data scales. RL overtakes the strongest SFT checkpoint (SFT-$16$k) after roughly $0.4$ M environment steps on OOD task. At convergence, it performs comparably to SFT-$16$k in the training setting and $42.6\,\%$ better on unseen objects and tables, showing that reinforcement learning not only improves policy performance under the training distribution but also provides markedly stronger generalization.
@@
We further evaluate the generalization performance of the converged RL policy and the best-performing SFT policy (SFT-$16$k, shortly denoted as SFT) on OOD tasks, as shown in Fig. 7. Here we report average success rates and average performance drop over three random seeds while full detailed results can be found in Sec. B.3. The relative performance drop is calculated as $P=\frac{\text{OOD}-\text{IND}}{\text{IND}}$. The results indicate that RL performs comparably to SFT in Vision tasks, noticeably better in Semantics, and significantly better in Execution.
@@ 5.4 Dissecting RL’s contributions to generalization
To probe qualitative differences between the two approaches, we visualise policy roll-outs on four representative tasks (Fig. 9). In Vision/Dynamic-Noise (strong), the SFT agent repeatedly drops the object immediately after grasping, indicating that heavy visual perturbations prevent it from localising the receptacle, whereas the RL agent completes the placement. With an Unseen Object in Semantics, SFT keeps attempting to grasp the item it already holds and stalls, while RL lifts it and finishes the task. This may be attributed to RL’s extensive trial-and-error experience during training, enabling it to make finer-grained control when interacting with unseen objects. RL also learned to recover from failed grasps and mid-episode object shifts in two Execution tests, whereas SFT marches on in spite of position errors, likely because such cases never appear in the demonstration data. The trajectory distributions encountered during training (Fig. 8) echo this gap: RL trajectories span a broader workspace and a richer range of end-effector orientations, whereas SFT roll-outs cluster along the motion-planner paths present in the dataset. This wider coverage appears key to RL’s superior generalisation on Execution tasks.
@@ A.1 Implementation of different fine-tuning algorithm
We follow the process supervision RL in GRPO (Shao et al., 2024), where the advantage is calculated as the summation of the normalized rewards from future steps
$\hat{A}^{i}_{t}$ $=\sum_{\text{index}(j)\geq t}\frac{r^{i}_{\text{index}(j)}-\text{mean}(\mathbf{R})}{\text{std}(\mathbf{R})}$ (3)
$,\displaystyle\mathbf{R}=\{\{r^{1}_{\text{index}(1)},$$\cdots,r^{1}_{\text{index}(K_{1})}\},\cdots,\{r^{G}_{\text{index}(1)},\cdots,r^{G}_{\text{index}(K_{G})}\}\}$ (4)
where $\text{index}(j)$ denotes the timestamp of the $j$-th valid (non-zero) reward, and $K_{i}$ is the total number of steps of the $i$-th sampled trajectory within a group.
Due to differences between robot tasks and natural language QA, we also experiment with an alternative implementation, where $\mathbf{R}_{w}$ is defined as the whole reward of the trajectory. In this case, all rewards, including zeros, are considered valid. Thus, we have
$\hat{A}^{i}_{t}$ $=\sum_{j\geq t}\frac{r^{i}_{j}-\text{mean}(\mathbf{R}_{w})}{\text{std}(\mathbf{R}_{w})}$ (5)
$,\displaystyle\mathbf{R}_{w}=\{\{r^{1}_{1},$$\cdots,r^{1}_{K_{1}}\},\cdots,\{r^{G}_{1},\cdots,r^{G}_{K_{G}}\}\}$ (6)
We compared the performance of different GRPO implementations, as reported in Fig. 10. “(s)” indicates that a group of trajectories is sampled from the same initial state, as described in Sec. 4.1. The results show that, regardless of the implementation, GRPO consistently underperforms compared to PPO, further highlighting PPO’s superior performance in robotic tasks.
Figure 10: Comparison between different implementation of GRPO.
@@
In the original TPO paper (Zhang et al., 2024b), the total reward consists of a task success reward, a self-evaluated score, and an LLM-generated dense reward. In our task, only a sparse reward is available, as described in Sec. 5.1. Therefore, we use only this sparse reward for simplicity.
@@ A.3 Details of assets usage
As described in Sec. 5.1, we acquire digital assets from multiple sources. The 3D assets of objects and receptacles are sourced from the ManiSkill simulator (Tao et al., 2024), Objaverse (Deitke et al., 2023), or downloaded from Sketchfab [^1] https://sketchfab.com/. Additional table appearances are synthesized using Stable Diffusion (Rombach et al., 2022) and ControlNet (Zhang et al., 2023a); specifically, we use the online service provided by LibLibAI [^2] https://liblib.art/. Distractor textures are downloaded from ambientCG [^3] https://ambientcg.com/. We comply with the licenses of these assets accordingly.

### v3

@@ 1 Introduction
To answer this, we center our investigations on the representative pick-and-place task. On this task, we first evaluate mainstream RL algorithms for large-scale models (PPO (Schulman et al., 2017; Ouyang et al., 2022), DPO (Rafailov et al., 2023; Zhang et al., 2024b), GRPO (Shao et al., 2024; Guo et al., 2025a)) to identify effective VLA fine-tuning strategies. We then conduct a broad comparative evaluation, also on pick-and-place, pinpointing where RL fine-tuning outperforms SFT. As previewed in Fig.˜1, this comprehensive evaluation rigorously examines generalization across three key dimensions: 1) Vision: challenging generalization with novel backgrounds (unseen tables) and by overlaying unseen textures on the foreground or the entire image. 2) Semantics: testing understanding via unseen objects, novel receptacles, and varied instruction phrasings to probe language sensitivity and object recognition. 3) Execution: probing robustness by varying initial robot states, object/receptacle positions, and introducing dynamic disturbances like random object reposition within episodes.
@@ 3.2 Vision-Language-Action models
We base our study on OpenVLA (Kim et al., 2024) (Fig.˜2), an open-source model that achieves state-of-the-art performance by pairing a fused visual encoder of SigLIP (Zhai et al., 2023) and DINOv2 (Oquab et al., 2023) with a Llama-2 7B language backbone (Touvron et al., 2023), built on the Prismatic VLM (Karamcheti et al., 2024). At each time step the policy receives a single RGB image $o_{t}$ and an instruction $l$, i.e., the history length $H=1$. The image is embedded into visual tokens, the instruction is tokenised with Llama 2’s tokenizer, and the resulting token sequence is fed to the causal transformer decoder.
@@ 4 Effective RL fine-tuning of VLA models
We begin our study by exploring how to fine-tune VLA models with reinforcement learning. As these models inherit a similar scale and structure from their pre-trained LLM/VLM backbones, we first test whether common RL algorithms for large-scale models like LLMs/VLMs transfer effectively to VLAs. In Sec.˜4.1, we will show that Proximal Policy Optimization (PPO) (Schulman et al., 2017) consistently improves performance, whereas Direct Preference Optimization (DPO) (Rafailov et al., 2023; Zhang et al., 2024b) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024) struggle to learn something for POMDP robotic tasks. Guided by this finding, we then ablate key design choices within PPO to distill an effective fine-tuning recipe for VLA models in Sec.˜4.2.
@@ 4.1 RL algorithms: PPO, GRPO, DPO
  • GRPO estimates the baseline using a group of samples, allowing for direct computation of advantages without explicit value estimation. While natural language generation and robotic multi-step MDP tasks are inherently different, we follow the GRPO setup (Shao et al., 2024) by sampling trajectories from a common initial state, with each group containing 8 trajectories.
  • DPO exploits offline datasets annotated with pairwise preferences. In robotic tasks, however, obtaining contrastive trajectories from the same initial state is difficult. Following (Zhang et al., 2024b), we infer trajectory-level preferences from reward signals. To ensure a fair comparison, we employ only the stage-based sparse reward described in Sec.˜5.1, rather than the richer reward scheme used in the original implementation by (Zhang et al., 2024b).
These algorithms are illustrated in Fig.˜3(a), with further details and implementation specifics provided in Sec.˜A.1. We evaluate the performance on the pick-and-place task described in Sec.˜5.1.
The results are presented in Fig.˜3(b), with each experiment conducted using two different random seeds. Our findings indicate that PPO consistently outperforms GRPO. We attribute this to fundamental differences in environment dynamics between natural language tasks and robotic tasks (Li et al., 2024b). We hypothesis that, in robotic tasks’ POMDP, each action sequentially alters the environment state in a non-stationary manner, which may destabilizes GRPO’s advantage estimates. Additionally, PPO surpasses DPO performance. We hypothesize that this is due to the sparse reward structure, which makes it challenging to distinguish between the quality of different trajectories, as well as significant distribution shifts between the offline dataset and interactive execution (Prudencio et al., 2023).
@@ 4.2 Design factors of PPO
Since PPO—and most modern RL methods—use an actor–critic formulation (Konda and Tsitsiklis, 1999), we treat the pretrained VLA policy as the actor and attach a lightweight critic that estimates the state value $V(s)$. To keep the architecture compact, the actor and critic share the entire Transformer backbone; a three-layer MLP value head (see Fig.˜4(a)) takes the hidden vector $h^{0}$ produced by the final Transformer block (Vaswani et al., 2017) at the first action-token position and regresses it to a scalar value. To verify its efficacy, we evaluated several critic designs. Feeding the value head with the first action-token embedding $h^{0}$ produced the highest and most stable returns, outperforming the last-token input $h^{n}$ and the concatenation $h^{0},\dots,h^{n}]$ (see Fig.˜4(b)). A critic that uses its own Transformer backbone achieved comparable rewards but trained $35\,\%$ slower and consumed $83\,\%$ more VRAM ($81.3$ GB vs. $44.4$ GB). These results confirm that a shared backbone with $h^{0}$ is the most efficient choice.
@@
We run all experiments with the official OpenVLA checkpoint (Kim et al., 2024) pretrained on the OXE dataset (Collaboration et al., 2023). Due to the poor performance of this checkpoint on our benchmark, we warm it up with $140$ demonstration trajectories gathered by Octo-Small (Team et al., 2024) and a motion planner (see Sec.˜A.2). As shown in Fig.˜5(a), the warm-up model reaches convergence with roughly $50\,\%$ fewer environment steps, while both initialisations attain comparable asymptotic returns given sufficient interaction. Unless stated otherwise, all subsequent experiments use the warmed-up OpenVLA model for initialization.
@@
We show that the update–to–data ratio is a key factor for efficient fine-tuning. In PPO this ratio is set by the epoch hyper-parameter, which specifies how many gradient passes each batch receives. Fig.˜5(b) and Fig.˜5(c) reveal that increasing the PPO epoch beyond one yields no gains in return or sample efficiency, yet lengthens wall-clock time almost linearly. Accordingly, we fix $\text{epoch}=1$ in all remaining experiments, achieving the fastest training without sacrificing performance.
@@ 5.1 Environments and datasets
To thoroughly investigate the generalization capabilities of VLA, we focus on a typical pick-and-place task (Kim et al., 2025; Li et al., 2024a), where the agent is instructed to place an object from the table into a receptacle. Inspired by prior works (Fan et al., 2025; Stone et al., 2023) and the concept of Vision-Language-Action models, we define three dimensions of generalization: Vision, Semantics (Language), and Execution (Action), as illustrated in Fig.˜1.
@@
To probe generalisation, we randomise each task along three axes during training: Vision (16 tables), Semantics (16 objects), and Execution (perturbations of object and receptacle poses). At test time we hold at least one of these factors out of distribution, introducing nine novel objects, sixteen unseen receptacles, five new table surroudings, and sixteen distractor textures. Assets are drawn from Objaverse (Deitke et al., 2023) and other public sources; additional table appearances are synthesised with Stable Diffusion (Rombach et al., 2022) and ControlNet (Zhang et al., 2023a), ensuring surface variation without changing the table’s global pose. More details of asset acquisition and task specification are provided in Secs.˜A.3 and A.4. Built upon these assets, our tasks run in ManiSkill (Tao et al., 2024) with an 8-DoF WidowX-250S arm. At every step the agent observes a $640\times 480$ RGB frame and a natural-language instruction, and outputs a Cartesian end-effector delta plus a binary gripper signal. Rewards are sparse: $0.1$ for grasping and continuously holding the correct object, and $1.0$ for placing it successfully. For supervised fine-tuning we collect demonstration trajectories using the MPLib motion planner (Guo et al., 2024) and fine-tuned using LoRA (Hu et al., 2022); further dataset details can be found in Sec.˜A.2.
@@ 5.2 Effect of data scale on SFT performance
Following Lin et al. (2025), we study how supervised fine-tuning (SFT) scales with demonstration count. We train OpenVLA to convergence on datasets ranging from a few hundred to 6464k expert trajectories ($\sim 1.26$M transitions) and report average scores over three random seeds in-distribution and on unseen objects/tables (Fig.˜6(a) and 6(b)). We observe that performance plateaus at roughly 16k trajectories in both settings, so we adopt the 16k-trajectory SFT checkpoint as the baseline for comparing RL fine-tuning methods.
@@ 5.3 Performance comparison between RL and SFT
Building on preceding analyses, we plot success rates during RL training for both the training distribution and the OOD object/table split in Fig.˜6(c) and 6(d), alongside SFT scores at various data scales. RL overtakes the strongest SFT checkpoint (SFT-1616k) after roughly $0.4$ M environment steps on OOD task. At convergence, it performs comparably to SFT-1616k in the training setting and $42.6\,\%$ better on unseen objects and tables, showing that reinforcement learning not only improves policy performance under the training distribution but also provides markedly stronger generalization.
@@
We further evaluate the generalization performance of the converged RL policy and the best-performing SFT policy (SFT-1616k, shortly denoted as SFT) on OOD tasks, as shown in Fig.˜7. Here we report average success rates and average performance drop over three random seeds while full detailed results can be found in Sec.˜B.3. The relative performance drop is calculated as $P=\frac{\text{OOD}-\text{IND}}{\text{IND}}$. The results indicate that RL performs comparably to SFT in Vision tasks, noticeably better in Semantics, and significantly better in Execution.
@@ 5.4 Dissecting RL’s contributions to generalization
To probe qualitative differences between the two approaches, we visualise policy roll-outs on four representative tasks (Fig.˜9). In Vision/Dynamic-Noise (strong), the SFT agent repeatedly drops the object immediately after grasping, indicating that heavy visual perturbations prevent it from localising the receptacle, whereas the RL agent completes the placement. With an Unseen Object in Semantics, SFT keeps attempting to grasp the item it already holds and stalls, while RL lifts it and finishes the task. This may be attributed to RL’s extensive trial-and-error experience during training, enabling it to make finer-grained control when interacting with unseen objects. RL also learned to recover from failed grasps and mid-episode object shifts in two Execution tests, whereas SFT marches on in spite of position errors, likely because such cases never appear in the demonstration data. The trajectory distributions encountered during training (Fig.˜8) echo this gap: RL trajectories span a broader workspace and a richer range of end-effector orientations, whereas SFT roll-outs cluster along the motion-planner paths present in the dataset. This wider coverage appears key to RL’s superior generalisation on Execution tasks.
@@ 6 Conclusion
Acknowledgements
This research was supported by National Natural Science Foundation of China (No.62325405,62406159), Tsinghua University Initiative Scientific Research Program, Tsinghua-Efort Joint Research Center for EAI Computation and Perception, Beijing National Research Center for Information Science, Technology (BNRist), Beijing Innovation Center for Future Chips, and State Key laboratory of Space Network and Communications.
We would like to express our sincere gratitude to Liangzhi Shi, who carried out the deployment on the Franka Panda robotic arm and provided valuable sim-to-real evaluation results. We are also grateful to Hongzhi Zang for her insightful advice on GRPO performance tuning, which significantly improved the quality of this work.
@@ A.1 Implementation of different fine-tuning algorithm
We follow the outcome supervision RL paradigm in GRPO (Shao et al., 2024), where the advantage is computed as the normalized outcome rewards within each group:
$\hat{A}^{i}_{t}$ $=\frac{r^{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}$ (3)
$\mathbf{r}$ $=\{r^{1},r^{2},\cdots,r^{G}\}$ (4)
where $r^{i}$ is the outcome reward for trajectory $i$, and $G$ denotes the number of trajectories in a group.
To stabilize training, we use a total of 32 groups, each containing 8 trajectories. For successful placement episodes, the trajectory is truncated immediately after the first successful step. In contrast, unsuccessful episodes remain untruncated. The outcome reward is given only at the final step of each episode and is defined as the sum of the grasping, holding, and placement rewards.
@@
In the original TPO paper (Zhang et al., 2024b), the total reward consists of a task success reward, a self-evaluated score, and an LLM-generated dense reward. In our task, only a sparse reward is available, as described in Sec.˜5.1. Therefore, we use only this sparse reward for simplicity.
@@ A.3 Details of assets usage
As described in Sec.˜5.1, we acquire digital assets from multiple sources. The 3D assets of objects and receptacles are sourced from the ManiSkill simulator (Tao et al., 2024), Objaverse (Deitke et al., 2023), or downloaded from Sketchfab [^1] https://sketchfab.com/. Additional table appearances are synthesized using Stable Diffusion (Rombach et al., 2022) and ControlNet (Zhang et al., 2023a); specifically, we use the online service provided by LibLibAI [^2] https://liblib.art/. Distractor textures are downloaded from ambientCG [^3] https://ambientcg.com/. We comply with the licenses of these assets accordingly.